# Gauge freedoms in unravelled quantum dynamics: When do different continuous measurements yield identical quantum trajectories?

### v1

@@ VII.2 Proof of Theorem 3
for which taking $f(\psi,\bm{Q})=g(\psi)$ we recover again Eq. (51).
@@
$0=\left[\tilde{\mathcal{B}}(\psi)-{\cal B}(\psi)\right]\cdot\nabla f(\psi,\bm{Q})\\ +\sum_{\alpha}\int d\psi^{\prime}\Big{\{}\tilde{W}_{\alpha}(\psi,\psi^{\prime})\big{[}f(\psi^{\prime},\bm{Q}+\bm{E}_{\pi(\alpha)})-f(\psi,\bm{Q})\big{]}\\ -W_{\alpha}(\psi,\psi^{\prime})\big{[}f(\psi^{\prime},\bm{Q}+\bm{E}_{\alpha})-f(\psi,\bm{Q})\big{]}\Big{\}}$ (106)
@@
Combining this with Eq. (53), only the last term in Eq. (108) survives, and considering $\beta=\pi(\alpha)$ we obtain the partially-labelled jump condition,
$\tilde{W}_{\alpha}(\psi,\psi^{\prime})=W_{\pi(\alpha)}(\psi,\psi^{\prime})\quad\forall\alpha=1,\dots,d_{C},\quad\forall\psi,\psi^{\prime}\,.$ (109)
@@
$\delta\!\left\{\psi^{\prime}-\frac{\tilde{\mathcal{A}}_{\alpha}(\psi)}{\text{Tr}[\tilde{\mathcal{A}}_{\alpha}(\psi)]}\right\}\text{Tr}[\tilde{\mathcal{A}}_{\alpha}(\psi)]\\ =\delta\!\left\{\psi^{\prime}-\frac{\mathcal{A}_{\pi(\alpha)}(\psi)}{\text{Tr}[\mathcal{A}_{\pi(\alpha)}(\psi)]}\right\}\text{Tr}[\mathcal{A}_{\pi(\alpha)}(\psi)].$ (110)
@@
$\tilde{\mathcal{A}}_{\alpha}=\mathcal{A}_{\pi(\alpha)},\quad\forall\alpha=1,\dots,d_{C}\,.$ (111)
@@ VIII Discussion and outlook
These results provide a theoretical framework to investigate new phenomena in stochastic dynamics of open quantum systems. In particular, they are relevant when assessing whether a given phenomenon is robust to choosing different representations of the quantum master operator or continuous measurement schemes, or if its conditions are more restrictive. One relevant example is when a model exhibits weak unitary symmetry and the operation of the symmetry generates a new representation of the same QME. This symmetry remains at the level of the unravelled quantum dynamics when the symmetry transformed representation has a generator which is the same as that for the initial representation. This is studied in [47] which exploits directly the theorems given in this paper. The results presented here could also be useful for study of non-unitary symmetries exhibited by open quantum systems [66, 67, 68, 69].
@@ A.1 Types of SJEDs
Multiplying from the left by state vectors orthogonal to $J_{k^{\prime}}|\psi_{1}\rangle$ and $J_{k^{\prime}}|\psi_{2}\rangle$, this can be satisfied in two ways: either $c_{1}=c_{2}=c_{3}$, or $|\psi_{1}\rangle$ and $|\psi_{2}\rangle$ are parallel such that $|\psi_{2}\rangle=c_{4}|\psi_{1}\rangle$ with $c_{3}=(c_{1}+c_{2}c_{4})/(1+c_{4})$. Fixing $|\psi_{1}\rangle$, these results must hold for all $|\psi_{2}\rangle$ with $J_{k}|\psi_{2}\rangle\neq 0$. If $c_{2}=c_{1}$ for all $|\psi_{2}\rangle$, this means that $J_{k^{\prime}}=c_{1}J_{k}$ and the SJED is of non-reset type. Otherwise $J_{k^{\prime}}|\psi_{2}\rangle$ must be always parallel to $J_{k}|\psi_{2}\rangle$, which means that $J_{k},J_{k^{\prime}}$ both have rank 1, and the SJED is of reset type.
@@ A.3 Combined Representations of SJEDs
We now consider situations with more than one SJED. As in Sec. III.3 we suppose that $H^{\prime},J_{1}^{\prime},\dots,J_{d^{\prime}}^{\prime}$ is a representation of the QME in which all SJEDs have minimal representations. There are $d_{C}$ SJEDs which are $S^{\prime}_{1},\dots,S^{\prime}_{d_{C}}$. Let ${H},J_{1},\dots,J_{d}$ be a generic representation with SJEDs $S_{1},\dots,S_{d_{C}}$. We will show that conditions Eq. (28) of Theorem 1 are equivalent to the conditions in Eq. (31).
@@
To do this in a systematic way, we start with an arbitrary partitioning of the indices $1,\dots,d$ into sets ${S}_{1},\dots,{S}_{d_{C}}$. Choose an arbitrary permutation $\pi$ of $\{1,\dots,d_{C}\}$. Then jump operator $J_{j}$ can be obtained as
${J}_{j}=\sum_{k\in S^{\prime}_{\pi(\alpha)}}\hat{\mathbf{V}}^{(\alpha)}_{jk}J_{k}^{\prime}\quad\text{for}\quad j\in{S}_{\alpha},$ (122)
where $\hat{\mathbf{V}}^{(\alpha)}$ is similar to the isometry from Eqs. (118) and (121), except that it is now defined for $k\in S^{\prime}_{\pi(\alpha)}$ [instead of $k\in S^{\prime}_{\alpha}$], to allow for permutation of the SJEDs. The relevant isometric property becomes
$\sum_{j\in S_{\alpha}}\left[\hat{\mathbf{V}}^{(\alpha)}_{jk}\right]^{*}\hat{\mathbf{V}}^{(\alpha)}_{jk^{\prime}}=\delta_{kk^{\prime}}\quad\text{for}\quad k,k^{\prime}\in S^{\prime}_{\pi(\alpha)}\;.$ (123)
@@
$\mathbf{V}_{jk}^{(\alpha)}=\begin{cases}\hat{\mathbf{V}}_{jk}^{(\alpha)}\quad&\text{if}\quad j\in S_{\alpha},\,k\in S^{\prime}_{\pi(\alpha)}\\ 0&\text{otherwise}\;.\end{cases}$ (124)
@@
${J}_{j}=\sum_{k\in S^{\prime}_{\pi(\alpha)}}\mathbf{V}^{(\alpha)}_{jk}J_{k}^{\prime}\quad\text{for}\quad j\in{S}_{\alpha},$ (125)
@@ Appendix B Additional Example
We discuss an example in which the permutation appearing in Theorems 1 and 3 is non-trivial. As in Sec. V, we consider a 3-level system with the basis $|0\rangle$, $|1\rangle$, $|2\rangle$ and with an arbitrary Hamiltonian $H$. Let the jump operators be
@@
$\theta=0^{\circ}:\quad$ $\tilde{\cal A}_{1}={\cal A}_{1},\qquad\tilde{\cal A}_{2}={\cal A}_{2},$
$\theta=90^{\circ}:\quad$ $\tilde{\cal A}_{1}={\cal A}_{2},\qquad\tilde{\cal A}_{2}={\cal A}_{1}.$ (131)
@@ Appendix C Pure state with distinct destinations for all SJEDs
Here, we describe how to find $\psi_{0}$ which is at the center of $\mathbf{C}$ in Eq. (62) that is used in the proof of Theorem 1. This discussion refers back to the properties (i)-(iii) of set $\mathbf{C}$, given at the beginning of Sec. VI.2.1.
Start from an arbitrary candidate state $\psi=|\psi\rangle\!\langle\psi|$. (In the following, we sometimes denote states via matrices $\psi$ and sometimes via the corresponding vectors $|\psi\rangle$). Suppose that $\psi$ violates the conditions above because of two “degeneracies”, that $\mathcal{D}_{k}(\psi)=0$ and $\mathcal{D}_{k^{\prime}}(\psi)=\mathcal{D}_{k^{\prime\prime}}(\psi)\neq 0$ for some $k,k^{\prime},k^{\prime\prime}$ (all in different SJEDs). (These violate conditions (i) and (ii) respectively.) Then choose some $|\varphi\rangle$ such that $J_{k}|\varphi\rangle\neq 0$ and write $|\psi^{\prime}\rangle=(|\psi\rangle+a|\varphi\rangle)/z$ with $a\in\mathbb{R}$, where $z$ is a normalisation constant. Then $\mathcal{D}_{k}(\psi^{\prime})\neq 0$, so condition (i) is now satisfied. Moreover, taking sufficiently small $a>0$ ensures that this replacement does not generate any new degeneracies [for example, it avoids the situation that some $\mathcal{D}_{k^{\prime\prime\prime}}(\psi)\neq 0$ but $\mathcal{D}_{k^{\prime\prime\prime}}(\psi^{\prime})=0$].

### v3

@@ 7.2 Proof of Theorem 3
and taking $f(\psi,\bm{Q})=g(\psi)$ we recover again Eq. (51).
@@
$0=\left[\tilde{\mathcal{B}}(\psi)-{\cal B}(\psi)\right]\cdot\nabla f(\psi,\bm{Q})\\ +\sum_{\alpha}\int d\psi^{\prime}\Big{\{}\tilde{W}_{\alpha}(\psi,\psi^{\prime})\big{[}f(\psi^{\prime},\bm{Q}+\bm{E}_{\pi_{\textsc{c}}(\alpha)})-f(\psi,\bm{Q})\big{]}\\ -W_{\alpha}(\psi,\psi^{\prime})\big{[}f(\psi^{\prime},\bm{Q}+\bm{E}_{\alpha})-f(\psi,\bm{Q})\big{]}\Big{\}}$ (106)
@@
Combining this with Eq. (53), only the last term in Eq. (108) survives, and considering $\beta=\pi_{\textsc{c}}(\alpha)$ we obtain the partially-labelled jump condition,
$\tilde{W}_{\alpha}(\psi,\psi^{\prime})=W_{\pi_{\textsc{c}}(\alpha)}(\psi,\psi^{\prime})\quad\forall\alpha=1,\dots,d_{C},\quad\forall\psi,\psi^{\prime}\,.$ (109)
@@
$\delta\!\left\{\psi^{\prime}-\frac{\tilde{\mathcal{A}}_{\alpha}(\psi)}{\text{Tr}[\tilde{\mathcal{A}}_{\alpha}(\psi)]}\right\}\text{Tr}[\tilde{\mathcal{A}}_{\alpha}(\psi)]=\delta\!\left\{\psi^{\prime}-\frac{\mathcal{A}_{\pi_{\textsc{c}}(\alpha)}(\psi)}{\text{Tr}[\mathcal{A}_{\pi_{\textsc{c}}(\alpha)}(\psi)]}\right\}\text{Tr}[\mathcal{A}_{\pi_{\textsc{c}}(\alpha)}(\psi)].$ (110)
@@
$\tilde{\mathcal{A}}_{\alpha}=\mathcal{A}_{\pi_{\textsc{c}}(\alpha)},\quad\forall\alpha=1,\dots,d_{C}\,.$ (111)
@@ 8 Discussion and outlook
These results provide a theoretical framework to investigate new phenomena in stochastic dynamics of open quantum systems. In particular, they are relevant when assessing whether a given phenomenon is robust to choosing different representations of the quantum master operator or continuous measurement schemes, or if its conditions are more restrictive. Any thermodynamic description that depends only on properties of quantum trajectories is naturally invariant under the gauge transformations discussed here [69, 23], compare [24]. Similarly, any dynamical parameters encoded by gauge transformations of unravelled dynamics cannot be inferred from quantum trajectories or coarse-grained measurement records, instead requiring full measurement records [70, 71, 72, 73]. Another relevant example is when a model exhibits weak unitary symmetry and the operation of the symmetry generates a new representation of the same QME. This symmetry remains at the level of the unravelled quantum dynamics when the symmetry transformed representation has a generator which is the same as that for the initial representation. This is studied in [52] which exploits directly the theorems given in this paper. The results presented here could also be useful for study of non-unitary symmetries exhibited by open quantum systems [74, 75, 76, 77].
This work was supported by the Engineering and Physical Sciences Research Council [grant number EP/T517847/1].
@@ A.1 Types of SJEDs
Multiplying, separately, from the left by state vectors orthogonal to $J_{k^{\prime}}|\psi_{1}\rangle$ and $J_{k^{\prime}}|\psi_{2}\rangle$, this can be satisfied in two ways: either $c_{1}=c_{2}=c_{3}$, or $J_{k^{\prime}}|\psi_{1}\rangle$ and $J_{k^{\prime}}|\psi_{2}\rangle$ are parallel such that $J_{k^{\prime}}|\psi_{2}\rangle=c_{4}J_{k^{\prime}}|\psi_{1}\rangle$ with $c_{3}=(c_{1}+c_{2}c_{4})/(1+c_{4})$. Fixing $|\psi_{1}\rangle$, these results must hold for all $|\psi_{2}\rangle$ with $J_{k}|\psi_{2}\rangle\neq 0$. If $c_{1}=c_{2}$ for all $|\psi_{2}\rangle$, then from (112) we have that $J_{k}=c_{1}J_{k^{\prime}}$. Otherwise $J_{k^{\prime}}|\psi_{1}\rangle$ must be always parallel to $J_{k^{\prime}}|\psi_{2}\rangle$ (which in turn is parallel to $J_{k}|\psi_{2}\rangle$), therefore $J_{k},J_{k^{\prime}}$ both have rank 1, and the SJED is of reset type. Note from the definition of SJED types that we choose $J_{k^{\prime}}|\psi_{1}\rangle$ and $J_{k}^{\prime}|\psi_{2}\rangle$ being parallel to take precedence over $c_{1}=c_{2}$, so that SJEDs contain either only reset or non-reset jumps, according to their type.
@@ A.3 Combined Representations of SJEDs
We now consider situations with more than one SJED. As in Sec. 3.3 we suppose that $H^{\prime},J_{1}^{\prime},\dots,J_{d^{\prime}}^{\prime}$ is a representation of the QME in which all SJEDs have minimal representations. There are $d_{C}$ SJEDs which are $S^{\prime}_{1},\dots,S^{\prime}_{d_{C}}$. Let ${H},J_{1},\dots,J_{d}$ be a generic representation with SJEDs $S_{1},\dots,S_{d_{C}}$. We will show that conditions Eq. (28) of Theorem 1 are equivalent to the conditions in Eq. (31).
@@
To do this in a systematic way, we start with an arbitrary partitioning of the indices $1,\dots,d$ into sets ${S}_{1},\dots,{S}_{d_{C}}$. Choose an arbitrary permutation $\pi_{\textsc{c}}$ of $\{1,\dots,d_{C}\}$. Then jump operator $J_{j}$ can be obtained as
${J}_{j}=\sum_{k\in S^{\prime}_{\pi_{\textsc{c}}(\alpha)}}\hat{\mathbf{V}}^{(\alpha)}_{jk}J_{k}^{\prime}\quad\text{for}\quad j\in{S}_{\alpha},$ (122)
where $\hat{\mathbf{V}}^{(\alpha)}$ is similar to the isometry from Eqs. (118) and (121), except that it is now defined for $k\in S^{\prime}_{\pi_{\textsc{c}}(\alpha)}$ [instead of $k\in S^{\prime}_{\alpha}$], to allow for permutation of the SJEDs. The relevant isometric property becomes
$\sum_{j\in S_{\alpha}}\left[\hat{\mathbf{V}}^{(\alpha)}_{jk}\right]^{*}\hat{\mathbf{V}}^{(\alpha)}_{jk^{\prime}}=\delta_{kk^{\prime}}\quad\text{for}\quad k,k^{\prime}\in S^{\prime}_{\pi_{\textsc{c}}(\alpha)}\;.$ (123)
@@
$\mathbf{V}_{jk}^{(\alpha)}=\begin{cases}\hat{\mathbf{V}}_{jk}^{(\alpha)}\quad&\text{if}\quad j\in S_{\alpha},\,k\in S^{\prime}_{\pi_{\textsc{c}}(\alpha)}\\ 0&\text{otherwise}\;.\end{cases}$ (124)
@@
${J}_{j}=\sum_{k\in S^{\prime}_{\pi_{\textsc{c}}(\alpha)}}\mathbf{V}^{(\alpha)}_{jk}J_{k}^{\prime}\quad\text{for}\quad j\in{S}_{\alpha},$ (125)
@@ Appendix B Additional Example
We discuss an example in which the permutation appearing in Theorems 1 and 3 is non-trivial. As in Sec. 5, we consider a 3-level system with the basis $|0\rangle$, $|1\rangle$, $|2\rangle$ and with an arbitrary Hamiltonian $H$. Let the jump operators be
@@
$\theta=0^{\circ}:$ $\tilde{\cal A}_{1}={\cal A}_{1},\qquad\tilde{\cal A}_{2}={\cal A}_{2},$
$\theta=90^{\circ}:$ $\tilde{\cal A}_{1}={\cal A}_{2},\qquad\tilde{\cal A}_{2}={\cal A}_{1}.$ (131)
@@ Appendix C Pure state with distinct destinations for all SJEDs
Here, we describe how to find $\psi_{0}$ which is at the center of $\mathbf{C}$ in Eq. (62) that is used in the proof of Theorem 1. This discussion refers back to the properties (i)-(iii) of set $\mathbf{C}$, given at the beginning of Sec. 6.2.1.
Start from an arbitrary candidate state $\psi=|\psi\rangle\!\langle\psi|$. (In the following, we sometimes denote states via matrices $\psi$ and sometimes via the corresponding vectors $|\psi\rangle$). Suppose that $\psi$ violates the conditions (i,ii) above, because of two “degeneracies”, that $\mathcal{D}_{k}(\psi)=0$ and $\mathcal{D}_{k^{\prime}}(\psi)=\mathcal{D}_{k^{\prime\prime}}(\psi)\neq 0$ for some $k,k^{\prime},k^{\prime\prime}$ (all in different SJEDs). Then choose some $|\varphi\rangle$ such that $J_{k}|\varphi\rangle\neq 0$ and write $|\psi^{\prime}\rangle=(|\psi\rangle+a|\varphi\rangle)/z$ with $a\in\mathbb{R}$, where $z$ is a normalisation constant. Then $\mathcal{D}_{k}(\psi^{\prime})\neq 0$, so condition (i) is now satisfied. Moreover, taking sufficiently small $a>0$ ensures that this replacement does not generate any new degeneracies [for example, it avoids the situation that some $\mathcal{D}_{k^{\prime\prime\prime}}(\psi)\neq 0$ but $\mathcal{D}_{k^{\prime\prime\prime}}(\psi^{\prime})=0$].